OpenStudy (anonymous):

Solve using Quadratic Formula: 2x^-2-x^-1-1=O

OpenStudy (anonymous):

\[x^{-2} =\frac{1}{x^2}\]

OpenStudy (anonymous):

can u show me ur solution?

OpenStudy (anonymous):

can you please check your equation, is it is -2-x^-1???

OpenStudy (anonymous):

|dw:1316399903401:dw

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[2x^-2-x^ -1-1=O\]

OpenStudy (anonymous):

the exponents are negative

OpenStudy (anonymous):

so it will be: 2/x^2 - 1/x - 1 = 0 2/x^2 - 1/x = 1 ----> multiply both side by x^2 2 - x = x^2 x^2 + x - 2 = 0 (x+2)(x-1) = 0 x = -2 x = 1 :-)

OpenStudy (anonymous):

just make the negative exponential variable upside down( take the reciprocal) to make the exponent positive. :-)

OpenStudy (anonymous):

basta pag may hindi ka maintindihan ask ka lng. :-)

OpenStudy (anonymous):

ok, thanks:)